# Boosting Robustness by Clipping Gradients in Distributed Learning

### v1

@@ 4 Adaptive Robust Clipping (ARC) and its Properties
In this section, upon presenting a preliminary observation on the fragility of static clipping, we introduce adaptive robust clipping (i.e., ARC). Lastly, we present some useful properties of ARC.
@@
This means that if the clipping threshold is a priori fixed, i.e., independent from the input vectors, pre-aggregation clipping does not preserve the robustness of the original aggregation. This fragility of such static clipping is also apparent in practice, as shown by our experimental study in Appendix F.
Description of ARC. In ARC, the clipping threshold is determined according to the input vectors. More precisely, ARC clips the largest $k=\left\lfloor{2(f/n)(n-f)}\right\rfloor$ vectors using a clipping parameter given by the norm of the $(k+1)$-th largest input vector. The overall scheme is formally presented in Algorithm 2, and its computational complexity is $\mathcal{O}(nd+n\log(n))$ (see Appendix A). ARC is adaptive in the sense that the clipping threshold $C$ is not fixed but depends on the input vectors.
@@
Properties of ARC. We present below important properties of ARC. Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation rule and $\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})\coloneqq\mathbf{F}(\mathbf{ARC}(x_{1},\dots,x_{n}))$
@@
Since $\kappa\geq\frac{f}{n-2f}$ (recalled in Section 3), Theorem 4.2 implies that $\mathbf{F}\circ\mathbf{ARC}$ is $\left(f,3\kappa\right)$-robust. In other words, ARC preserves the robustness of the original aggregation scheme.
Besides preserving robustness in the context of Robust-DGD, ARC also limits the norms of adversarial workers’ gradients by the norm of an honest worker’s gradient. Specifically, when $f/n<1/2$, the clipping threshold $C$ in ARC is bounded by $\max_{i\in\mathcal{H}}\left\lVert{x_{i}}\right\rVert$ (see Appendix C.2). This property is useful in proving our convergence result, stated below in Theorem 5.1. Proofs for Lemma 4.1 and Theorem 4.2 are deferred to Appendix B.
5 Learning Guarantee of Robust-DGD with ARC
In this section, we first present our convergence result for Robust-DGD with ARC (or Robust-DGD $\circ$ ARC), and then characterize its improvement over the lower bound recalled in Lemma 3.1. Specifically, throughout this section, we consider Algorithm 1 with aggregation $\mathbf{F}\circ\mathbf{ARC}$, i.e., in each learning step $t$, $R_{t}\coloneqq\mathbf{F}\circ\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\dots,\,{{g}^{\left({n}\right)}_{t}}\right)$ Note that, since we have already shown in Theorem 4.2 that $\mathbf{F}\circ\mathbf{ARC}$ is $(f,3\kappa)$-robust, a general convergence result for Robust-DGD $\circ$ ARC follows verbatim from Lemma 3.4, replacing $\kappa$ with $3\kappa$. However, we obtain below an alternative convergence result that enables us to show the improvement in learning error due to ARC.
5.1 Convergence of Robust-DGD with ARC
We assume that, for any set of vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$. This assumption can be made without loss of generality, the reason for which is deferred to Appendix C.1.
Theorem 5.1.
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\max\left\{5\kappa\left(G^{2}+B^{2}\rho^{2}\right),~{}\frac{2\Delta_{o}L}{T}+3\kappa\left(G^{2}+B^{2}\rho^{2}\right)\right\}.$
We have the following corollary of Theorem 5.1 and Lemma 3.4.
Corollary 5.2.
For parameters given in Theorem 5.1, if $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$, then Robust-DGD $\circ$ ARC achieves
$\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq 5\kappa~{}\min\left\{G^{2}+B^{2}\rho^{2},~{}\frac{G^{2}}{\max\{(1-\kappa B^{2}),0\}}\right\},$
5.2 Improvement in Robustness
In the following, we show that if the local gradients at the initial model ${{\theta}_{1}}$ are bounded, then Robust-DGD $\circ$ ARC overcomes the lower bound recalled in 3.1. Specifically, we make this assumption:
Assumption 5.3.
There exists $\zeta\in\mathbb{R}^{+}$ such that, at the initial model ${{\theta}_{1}}$, $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})}\right\rVert\leq\zeta$.
For ease of presentation, we denote $\mathsf{BP}\coloneqq\frac{1}{2+B^{2}},$
$\quad\varepsilon_{o}\left(f/n\right)\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$ $\Psi(G,B,\rho)\coloneqq 640\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right).$
Recall from Lemma 3.1 that $\mathsf{BP}$ and $\varepsilon_{o}(f/n)$ are the breakdown point and the achievable stationarity error when $f/n<\mathsf{BP}$, respectively, for the worst-case scenario under $(G,B)$-gradient dissimilarity. We obtain the following theorem for Robust-DGD $\circ$ ARC.
Theorem 5.4.
Suppose Assumption 5.3 and $B>0$. Let $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$, and let the remaining parameters for Robust-DGD $\circ$ ARC be as defined in Corollary 5.2. Let $\xi_{o}<1$ be a (strictly) positive real value, and let $\rho\coloneqq\exp\left(\frac{(2+B^{2})\Delta_{o}}{(1-\xi_{o})G^{2}}L\right)\zeta$. Then, the following holds true:
  1. 1.
    If $\frac{f}{n}\coloneqq(1-\xi)\mathsf{BP}$, where $\xi\in(0,\xi_{o}]$, then, $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\xi\,\Psi(G,B,\rho)\cdot\varepsilon_{o}\left(f/n\right)$
  2. 2.
    If $\mathsf{BP}\leq f/n<1/2$, then $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\Psi(G,B,\rho)\cdot\frac{G^{2}}{8}$
According to Theorem 5.4, if the local gradients at model initialization are bounded, then Robust-DGD $\circ$ ARC can tolerate a larger fraction of adversarial workers than the breakdown point $\mathsf{BP}$ under $(G,B)$-gradient dissimilarity. Moreover, even if $f/n<\mathsf{BP}$, Robust-DGD $\circ$ ARC can improve over the worst-case stationarity error under $(G,B)$-gradient dissimilarity, formally shown in the following corollary of Theorem 5.4. For the corollary, we consider the same parameters as in Theorem 5.4.
Corollary 5.5.
Suppose Assumption 5.3 and that $f/n\coloneqq(1-\xi)~{}\mathsf{BP}$, where $\xi\in(0,\xi_{o}]$. For $\upsilon\in(0,1)$, if $\xi\leq\frac{1}{\Psi(G,B,\rho)}\,\upsilon$, then $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\upsilon\cdot\varepsilon_{o}\left(f/n\right)$.
Corollary 5.5 implies that when local gradients at model initialization are bounded, Robust-DGD $\circ$ ARC improves over the lower bound $\varepsilon_{o}\left(f/n\right)$ for sufficiently small $\xi$ (or $f/n$ sufficiently close to $\mathsf{BP}$). [^5] An improvement over the lower bound automatically implies improvement over the original learning guarantee of Robust-DGD (recalled in Lemma 3.4), i.e., when model initialization is arbitrary. The closer $f/n$ is to $\mathsf{BP}$, the smaller is the factor of reduction $\upsilon$, and thus higher is the improvement. The threshold on $\xi$ approaches 00 when $B\rightarrow 0$, i.e., the space of improvement is quite narrow. This is expected, since the lower bound approaches $G^{2}f/(4(n-2f))$ in this case, which is already smaller than $5\kappa G^{2}$ (the error for Robust-DGD $\circ$ ARC in this case) for any $(f,\kappa)$-robust $\mathbf{F}$.
Reconciliation with the lower bound. When the workers’ gradients at model initialization can be arbitrarily large, the convergence guarantee of Robust-DGD $\circ$ ARC reduces to that of Robust-DGD. On the contrary, if the model initialization is good, i.e., $\zeta$ is small, then ARC improves the learning error, as soon as the tolerable fraction of adversarial workers is in the proximity of $\mathsf{BP}$. We remark that Assumption 5.3, pertaining to model initialization, is not too strong and can be satisfied in practice [15], also validated by our experimental study shown in Figure 17 of Appendix F. Effectively, under Assumption 5.3, ARC reduces the heterogeneity model to $(G+B\rho,0)$-gradient dissimilarity, with $\rho$ defined in Theorem 5.4, since $\max_{t\in[T]}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}({{\theta}_{t}})}\right\rVert\leq\rho$ in Robust-DGD $\circ$ ARC for the parameters defined in Theorem 5.1, shown implicitly in the proof of Theorem 5.1 in Appendix C. Proofs for all results presented in this section are deferred to Appendix C.
@@ 6 Empirical Evaluation
In this section, we delve into the practical performance of ARC when incorporated in Robust-DSGD (see Algorithm 3 in Appendix D), an order-optimal stochastic variant of Robust-DGD [3]. We conduct experiments on standard image classification tasks, covering different adversarial scenarios. We empirically test four SOTA aggregation rules when pre-composed with ARC. We also contrast these outcomes against the performance of Robust-DSGD when no gradient clipping is used. Our findings underscore that clipping workers’ gradients using ARC prior to the aggregation is crucial to ensure the robustness of existing algorithms, especially in extreme scenarios, i.e., when either the data heterogeneity is high or there is a large fraction of adversarial workers to be tolerated.
We plot in Table 1, and Figures 1 and 3, the metric of worst-case maximal accuracy. For each of the five Byzantine attacks executed (see Appendix D.3), we record the maximal accuracy achieved by Robust-DSGD during the learning procedure under that attack. The worst-case maximal accuracy is thus the smallest maximal accuracy reached across the five attacks. As the adversarial attack cannot be known in advance, this metric is critical to accurately evaluate the robustness of aggregation methods, since it provides an estimate of the potential worst-case performance of the algorithm. Furthermore, we use the Dirichlet [23] distribution of parameter $\alpha$ to simulate data heterogeneity in the workers’ datasets. The comprehensive experimental setup can be found in Appendix D. In this section, we present results on MNIST [11] and CIFAR-10 [31], and defer our findings on Fashion-MNIST [42] to Appendix E.2. Finally, for the sake of clarity, we mainly consider here the aggregations CWTM and GM, and show the remaining results on CWMed and MK in Appendix E.
@@
Figure 2: Performance of Robust-DSGD when using ARC and without clipping on MNIST. Among $n=11$, we consider $f=1$ adversarial worker executing FOE [43], and vary the heterogeneity.
ARC boosts robustness in high heterogeneity. Figure 2 shows the performance of Robust-DSGD against the FOE [43] attack when using ARC opposed to no clipping with $f=1$ adversarial worker among $n=11$.When the heterogeneity is low (i.e., $\alpha=0.5$, see Figure 5), the performance of ARC and no clipping are comparable for both aggregations. When the heterogeneity increases ($\alpha=0.1$), the benefits of using ARC are more visible. CWTM $\circ$ NNM and GM $\circ$ NNM significantly struggle to learn (with a very large variance), while the same aggregations composed with ARC almost match the accuracy of DSGD towards the end of the training. In extreme heterogeneity, the improvement induced by our method is the most pronounced, as ARC enables both aggregations to reach a final accuracy close to 90%. Contrastingly, the same aggregations without clipping stagnate at an accuracy close to 10% throughout the training. Interestingly, only $f=1$ adversarial worker among $n=11$ (i.e., less than 10%) suffices to completely deteriorate the learning when no clipping is applied.
@@
Figure 3: Worst-case maximal accuracies achieved by Robust-DSGD, with and without ARC, on heterogeneously-distributed MNIST with $10$ honest workers. Left: $f=1$ adversarial worker among $n=11$ for varying levels of heterogeneity. Right: extreme heterogeneity for varying $f$.
@@
Varying heterogeneity for fixed $f$. Figure 3a compares the worst-case maximal accuracies achieved by Robust-DSGD when $f=1$, with ARC and without, for varying levels of heterogeneity. While CWTM $\circ$ NNM and GM $\circ$ NNM yield comparable accuracies with ARC in low heterogeneity ($\alpha\geq 0.5$), their performance significantly degrades when $\alpha$ drops below that threshold. Indeed, when $\alpha=0.1$, their accuracies drop below 65%, while ARC enables the same aggregations to maintain their accuracy at just below 98%. In extreme heterogeneity, the performance of the aggregations without clipping completely deteriorates with accuracies close to 15%. Contrastingly, ARC efficiently mitigates the Byzantine attacks, resulting in accuracies above 85% in the worst case for both aggregations. Similar plots for $f=3,5,7$ convey similar observations (see Appendix E.1). Essentially, in settings where gradient clipping is not needed (e.g., low heterogeneity or small $f$), ARC performs at least as well as no clipping. Nonetheless, ARC induces a significant improvement near the breakdown point of SOTA robust aggregation schemes (e.g., high heterogeneity or large $f$).
@@
CWTM $\circ$ NNM 51.6 $\pm$ 5.1 67.8 $\pm$ 0.9 40.7 $\pm$ 0.5 60.5 $\pm$ 1.2
GM $\circ$ NNM 41.2 $\pm$ 3.5 67.0 $\pm$ 1.0 16.0 $\pm$ 2.3 60.0 $\pm$ 2.0
CWMed $\circ$ NNM 43.4 $\pm$ 4.2 69.4 $\pm$ 0.8 13.7 $\pm$ 12.1 62.7 $\pm$ 1.2
MK $\circ$ NNM 50.9 $\pm$ 5.1 68.7 $\pm$ 0.5 40.5 $\pm$ 0.7 59.9 $\pm$ 1.9

### v3

@@ 3 Adaptive Robust Clipping (ARC) and its Properties
In this section, we first present a preliminary observation on the fragility of static clipping, and then introduce adaptive robust clipping (i.e., ARC) along with its robustness guarantees.
@@
Thus, if the clipping threshold is fixed, i.e., independent from the input vectors, pre-aggregation clipping does not preserve the robustness of the original aggregation. This fragility of such static clipping is also apparent in practice, as shown by our experimental study in Appendix F (and Figure 1).
Description of ARC. ARC is an adaptive clipping scheme that only makes use of the standard robustness parameter $\nicefrac{{f}}{{n}}$, i.e., the tolerable fraction of adversarial workers. ARC is adaptive in the sense that the clipping threshold $C$ is not fixed but depends on the input vectors. Specifically, ARC clips the largest $k=\left\lfloor{2(f/n)(n-f)}\right\rfloor$ vectors using a clipping parameter given by the norm of the $(k+1)$-th largest input vector. The overall scheme is formally presented in Algorithm 2, and its computational complexity is $\mathcal{O}(nd+n\log(n))$ (see Appendix A). Thus, pre-composing more computationally expensive schemes like NNM and multi-Krum, which have a complexity of $\mathcal{O}(dn^{2})$, with ARC does not introduce a significant overhead.
@@
Robustness Guarantee. We present below the preservation of robustness guaranteed by ARC. Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation rule and $\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})\coloneqq\mathbf{F}(\mathbf{ARC}(x_{1},\dots,x_{n}))$
@@
Proofs for the results presented in this section are deferred to Appendix B. Since $\kappa\geq\frac{f}{n-2f}$ (recalled in Section 2), Theorem 3.2 implies that $\mathbf{F}\circ\mathbf{ARC}$ is $\left(f,3\kappa\right)$-robust. In other words, ARC preserves the robustness of the original aggregation scheme. Therefore, a convergence result for Robust-DGD with ARC follows verbatim from Lemma 2.6, replacing $\kappa$ with $3\kappa$. Despite this multiplicative factor of 3, we observe in the next section that incorporating ARC consistently improves the empirical performance of classical aggregation methods. A more detailed theoretical explanation will be provided later in Section 5.1.
@@ 4 Empirical Evaluation
In this section, we delve into the practical performance of ARC when incorporated in Robust-DSGD (Algorithm 3 in Appendix D), an order-optimal stochastic variant of Robust-DGD (Allouah et al., 2023a). We conduct experiments on standard image classification tasks, covering different adversarial scenarios. We empirically test four aggregation methods when pre-composed with ARC. We also contrast these outcomes against the performance of Robust-DSGD when no gradient clipping is used. Our findings underscore that clipping workers’ gradients using ARC prior to the aggregation is crucial to ensure the robustness of existing algorithms, especially in extreme scenarios, i.e., when either the data heterogeneity is high or the fraction of adversarial workers is large.
@@
Figure 3: Performance of Robust-DSGD when using ARC and without clipping on MNIST. There are 10 honest workers and $f=1$ adversarial worker executing FOE (Xie et al., 2020).
We show in Table 1, and Figures 2a and 4, the metric of worst-case maximal accuracy. For each of the five Byzantine attacks executed (see Appendix D.3), we record the maximal accuracy achieved by Robust-DSGD during the learning procedure under that attack. The worst-case maximal accuracy is thus the smallest maximal accuracy reached across the five attacks. As the attack cannot be known in advance, this metric is critical to accurately evaluate the robustness of aggregation methods, since it provides an estimate of the potential worst-case performance of the algorithm. Furthermore, we use the Dirichlet (Hsu et al., 2019a) distribution of parameter $\alpha$ to simulate data heterogeneity. The comprehensive experimental setup can be found in Appendix D. In this section, we focus on presenting our results for the CWTM and GM aggregation methods applied to MNIST and CIFAR-10. Results for Fashion-MNIST, as well as for the CWMed and MK aggregations, are provided in Appendix E. All aggregation methods in our experiments are pre-composed with NNM (Allouah et al., 2023a), but we omit explicit mention of NNM in their names throughout the text for simplicity.
@@
Figure 4: Worst-case maximal accuracies achieved by Robust-DSGD, with and without ARC, on heterogeneously-distributed MNIST with $10$ honest workers. Left: $f=1$ adversarial worker among $n=11$ for varying levels of heterogeneity. Right: $\alpha=0.5$ for varying $f$.
ARC boosts robustness in high heterogeneity. Figure 3 shows the performance of Robust-DSGD against the FOE attack when using ARC opposed to no clipping. In low heterogeneity (i.e., $\alpha=0.5$, see Figure 7), the performances of ARC and no clipping are comparable for both aggregations. When the heterogeneity increases ($\alpha=0.1$), the benefits of using ARC are more visible. CWTM and GM significantly struggle to learn (with a very large variance), while the same aggregations composed with ARC almost match the accuracy of DSGD towards the end of the training. In extreme heterogeneity, the improvement induced by our method is the most pronounced, as ARC enables both aggregations to reach a final accuracy close to 90%. Contrastingly, the same aggregations without clipping stagnate at around 10% throughout the training. Interestingly, only $f=1$ adversarial worker among $n=11$ (i.e., less than 10%) suffices to completely deteriorate the learning when no clipping is applied, highlighting the necessity of ARC in extreme heterogeneity.
@@
Varying heterogeneity for fixed $f$. Figure 4a compares the worst-case maximal accuracies achieved by Robust-DSGD when $f=1$, with ARC and without, for varying levels of heterogeneity. While CWTM and GM yield comparable accuracies with ARC in low heterogeneity ($\alpha\geq 0.5$), their performance significantly degrades when $\alpha$ drops below that threshold. Indeed, when $\alpha=0.1$, their accuracies drop below 65%, while ARC enables the same aggregations to maintain their accuracy at just below 98%. In extreme heterogeneity, the performance of the aggregations without clipping completely deteriorates with accuracies close to 15%. Contrastingly, ARC efficiently mitigates the Byzantine attacks, resulting in accuracies above 85% in the worst case for both aggregations. Similar plots for $f=3,5,7$ convey similar observations (see Appendix E.1). Essentially, in low heterogeneity or when the fraction of adversarial workers is small, no clipping yields good empirical results and ARC performs at least as well as no clipping. However, ARC induces a significant improvement near the breakdown point of SOTA robust aggregation schemes (e.g., high heterogeneity or large $f$).
ARC increases the breakdown point in adverse settings. Figure 2a of Section 1 shows that for $f\in\{1,...,5\}$, Robust-DSGD with CWTM and GM completely fails to learn, consistently yielding worst-case maximal accuracies close to 15%. This suggests that $f=1$ constitutes the breakdown point for these aggregations in extreme heterogeneity. However, composing them with ARC increases the breakdown point of these aggregations to $f=3$. Indeed, for $f=1$ and 2, ARC enables Robust-DSGD to achieve accuracies greater than 85% in the worst case. However, when $f\geq 3$, the performance degrades, although CWTM with ARC is still able to reach a satisfactory accuracy close to 70%. Moreover, even when the heterogeneity is not large, ARC still produces a significant improvement when the fraction of adversarial workers increases in the system. Indeed, in Figure 4b , the performances of ARC and no clipping are comparable for $f\leq 3$. However, the improvement induced by ARC is much more visible when $f\geq 5$. Particularly when $f=7$, ARC enables CWTM and GM to reach accuracies greater than 80% in the worst case, whereas the same aggregations yield accuracies below 40% without clipping . This observation indicates the raise of the breakdown point due to ARC in practice. Plots for $\alpha=0.1$ and 1 convey the same observations in Appendix E.1.
Figure 5: FOE on heterogeneous CIFAR-10 with $\alpha=0.075$. $f=1$ among $n=17$.
Improved robustness on CIFAR-10. We also conduct experiments on the more challenging CIFAR-10 task with $n=17$ and $f=1$. Table 2 shows the worst-case maximal accuracies achieved by ARC and no clipping for four robust aggregations. For $\alpha=0.2$, ARC consistently outputs accuracies greater than 67% for all aggregations, while the same aggregations without clipping yield lower accuracies (with a larger variance across seeds). For instance, GM achieves 41.2% on average, i.e., 26% less than its counterpart with ARC. In the more heterogeneous setting $\alpha=0.075$, ARC enables all aggregations to reach worst-case maximal accuracies close to 60% (with a small variance). However, executing the same methods without clipping significantly deteriorates the performance of Robust-DSGD, with GM and CWMed achieving 16% and 13.7% in worst-case accuracy, respectively. This can also be seen in Figure 5, where FOE completely degrades the learning when ARC is not used.
@@
CWTM 51.6 $\pm$ 5.1 67.8 $\pm$ 0.9 40.7 $\pm$ 0.5 60.5 $\pm$ 1.2
GM 41.2 $\pm$ 3.5 67.0 $\pm$ 1.0 16.0 $\pm$ 2.3 60.0 $\pm$ 2.0
CWMed 43.4 $\pm$ 4.2 69.4 $\pm$ 0.8 13.7 $\pm$ 12.1 62.7 $\pm$ 1.2
MK 50.9 $\pm$ 5.1 68.7 $\pm$ 0.5 40.5 $\pm$ 0.7 59.9 $\pm$ 1.9